OpenStudy (lgg23):

The length of the curve by the equation xy=10, from x=a to x=b is given by L=integral of f(x)dx from a to b where f(x)=__?__

OpenStudy (inkyvoyd):

y=10/x given that x=/=0

OpenStudy (inkyvoyd):

so, f(x)=10/x, but x must not be 0 in the interval [a,b] I think.